[Ncep.nhc.nco_contacts] HSU Shift Log -- Technical Edition -- Sun Jun 25, 2017

HSU Shift Log nhc.tsbadmin at noaa.gov
Mon Jun 26 02:32:27 UTC 2017


HSU Shift Log for Sun Jun 25 2017 11:57:32 GMT-0400 (EDT)

Shift: A

Forecaster: E Blake

Technical Issues Summary:

ATCF issues:

N-AWIPS issues:

AWIPS-2/NCP issues:

Other Techincal issues: SHIPS-RI is still broken due to the lack of  
satellite input (STD Dev of the IR BR Temp). It is a notable loss because  
it seems like every RI forecast we have depends on it, and it is a little  
tough to gauge the possibility of Dora going faster than the other models  
without the predictor working.

Miscellaneous issues:
-------------- next part --------------
An HTML attachment was scrubbed...
URL: https://www.lstsrv.ncep.noaa.gov/pipermail/ncep.nhc.nco_contacts/attachments/20170626/9a484ee8/attachment.html 


More information about the Ncep.nhc.nco_contacts mailing list